# Talking to GDELT Through Knowledge Graphs

### v1

@@ 1 Introduction
In this work we study several approaches for communicating with a corpus of text via relevant text and knowledge graph (KG) representation and retrieval. The goal of this work is to gain an understanding of the benefits and drawbacks of these Retrieval Augmented Generation (RAG) approaches to studying a corpus of text when combined with a Large Language Model (LLM). In addition to this comparison analysis we also develop a new ontology for constructing KGs [8] from the Global Data on Events, Location, and Tone (GDELT) [^1] https://www.gdeltproject.org/ [13] dataset. We also demonstrate how to leverage a LLM to interact with these KGs to gain an understanding of their usefulness.
@@
GDELT describes its structure as a Global Knowledge Graph (GKG, specifically, we use the the Global Knowledge Graph edition 2 (GKG2) of GDELT). But in fact GDELT-GKG2 is implemented as multiple linked tables recording information about the relationship between articles and events, and thus effectively has the structure of a relational database. A level of effort is required to realize GKG2 properly in the mathematical form of a KG, effectively a graph database, derived from and consistent with its native relational database form. To facilitate this effort, we built an ontology as a typing schema for a graph database to mirror that of the relational database schema.
As a way to enhance LLM outputs, researchers and practitioners have been quick in applying LLMs to query and understand proprietary data through retrieval-augmented-generation (RAG)[14]. It has been shown that reasoning over the typical RAG framework, which only takes advantage of the unstructured text articles, fails to capture global information about the provided data [5, 26]. Motivated by this limitation, there has been recent interest in adapting these techniques to the case where our data has a graph structure, so that the LLM can directly ingest important relationships in the knowledge base [5, 7, 15, 30]. These techniques provide ways to automatically query and extract information stored in a KG without the user explicitly needing to understand query languages to probe their knowledge base. Typically, these AI-based search algorithms find subgraphs that can be used to answer a user-provided query.
@@ 2 Constructing a Knowledge Graph for GDELT
As previously mentioned, the GDELT-GKG2 dataset is not actually in the form of a knowledge graph, but rather a database of three related tables:
@@
The database schema for these three CSV files is shown in Fig. 1222See also [9].. The key characteristics of this database relational schema should be interpreted as follows:
@@
  • Fields in angled bracketsdelimited-$\langle\mbox{angled brackets}\rangle$⟨ angled brackets ⟩ indicate the field name in the schema.
@@ 3.1 GDELT Knowledge Graphs
[…]
@@ 3.3 Results
More nuanced queries might require combining the local and global approaches of the techniques used here. For example, neither the DKG nor vector store approaches could give a satisfactory information regarding Niki Fennoy. This may be due to the high specificity required by this query; the search capabilities may entirely miss information about a specific person. On the other hand, the DKG can be used to identify specific names, though often with unsatisfactory levels of information. For example, the node in the DKG corresponding to Niki Fennoy only connects to article nodes where they are mentioned. Interestingly, out of the 14 articles which are linked to Niki Fennoy in the GDELT dataset, only 3 of these mention the person in the associated text. The GraphRAG method provides good direction in finding this balance, placing individual entities on nodes and descriptive relationships on edges.
Despite individual flaws, combining vector store and knowledge graph techniques has the promise to boost information retrieval beyond either individual technique. Returning to the example of names, we can use the KG to filter the 201 articles we obtained on the Baltimore bridge collapse and search for articles referencing a given person. Using this filtered list, we can perform a refined vector search on those articles to search for further information.

### v3

@@ 1 Introduction
In this work we study several approaches for communicating with a corpus of text via relevant text and knowledge graph (KG) representation and retrieval facilitated by Large Language Models (LLMs). Our goal is to understand the benefits and drawbacks of Retrieval Augmented Generation (RAG) approaches to corpus management and anlysis when combined with an LLM. Throughout we use as a case study a novel KG derived from the Global Data on Events, Location, and Tone (GDELT) [^1] https://www.gdeltproject.org/ [13] dataset.
As a way to enhance LLM outputs, researchers and practitioners have been quick in applying LLMs to query and understand proprietary data through retrieval-augmented-generation (RAG) [14]. It has been shown that reasoning over the typical RAG framework, which only takes advantage of the unstructured text articles, fails to capture global information about the provided data [5, 27].
Motivated by this limitation, there has been recent interest in adapting these techniques to the case where our data has a graph structure, so that the LLM can directly ingest important relationships in the knowledge base [5, 7, 15, 31]. More specifically, KGs [8] are graph structures which are richly attributed with typing and semantic information on both nodes and edges. KG techniques provide ways to automatically query and extract information stored in a KG without the user explicitly needing to understand query languages to probe their knowledge base. Typically, these AI-based search algorithms find subgraphs that can be used to answer a user-provided query.
@@
GDELT describes its structure as a Global Knowledge Graph (GKG, specifically, we use the the Global Knowledge Graph edition 2 (GKG2) of GDELT). But in fact GDELT-GKG2 is implemented as multiple linked tables recording information about the relationship between articles and events, and thus effectively has the structure of a relational database. Another important contribution of this paper is to actually realize GKG2 properly in the mathematical form of a KG, effectively a graph database, derived from and consistent with its native relational database form. To facilitate this effort, we have identified a lightweight ontology for GDELT in the form of its graph schema, realizing its relational database schema in a KG form.
@@
The current state of neurosymbolic work is noted for the plethora of experimental architectures available. While details are explicated below in Section 3.2, we preview ours in Figure 1, including the five methodological pathways which are quantitatively compared: 1) graph queries on the KG (called the DKG) derived “directly” from GKG2; 2) use of G-Retriever [^2] https://github.com/XiaoxinHe/G-Retriever [7] against the same DKG; 3) RAG against a vector store representation of GKG2; 4) G-Retriever against a second KG (called the LKG) derived from using Llamaindex [^3] https://www.llamaindex.ai/ [1] against the GDELT source articles; and 5) GraphRAG [^4] https://microsoft.github.io/graphrag/ Q&A deployed against a third KG (called GRKG) using Microsoft’s open-source GraphRAG package with default configuration parameters.
@@ 2 Constructing a Knowledge Graph for GDELT
As previously mentioned, while the GDELT-GKG2 dataset is not actually natively in the form of a knowledge graph, it is advertised and frequently cited as being one. We believe that we are making a distinct contribution to the research community by converting the very popular GKG2 database into a proper KG.
GKG2 is natively a database of three related tables:
@@
The database schema for these three CSV files is shown in Fig. 2 (see also [9]). The key characteristics of this database relational schema should be interpreted as follows:
@@
Methods to automatically determine the graphical form of a relational database are widely known [21]. Most naturally, consider a table $T$ with $m$ rows $T[j],1\leq j\leq m$ and $n$ columns $T.i,1\leq i\leq n$. Then each of the $m$ rows $T[j]$ is represented as a node in one meta-class labeled by the primary keys. This node then has $n$ outgoing edges, each connecting to a node in another meta-class representing the field value $T[j].i$, and labeled by the column name. The resulting “star” bipartite graphs are then linked over shared values, including across multiple tables. This method straightforwardly produces a graph schema consistent with a given RDB, which may or may not be of sufficient complexity to warrant the lofty description of “ontology”. In our case, such a straightforward approach is mostly warranted, although as we will see additional constraints in the event table will argue for a somewhat more specific and idosyncratic graph structure.
@@
[…]
@@ 3.1 GDELT Knowledge Graphs
(c) Example GRKG constructed corpus of text using GraphRAG, removing all isolated nodes. Large nodes have degree $\geq 25$.
@@ 3.3 Results
To quantitatively evaluate the quality of the answers generated by our different question-answering methods, we require a set of ground truth answers for the questions posed. Table 2 presents these manually curated ground truth answers, representing the expected correct responses for each question used in our evaluation. These ground truth answers serve as the benchmark against which the performance of each system is measured.
Question Ground Truth
What is the name of the Bridge that collapsed and what river was it on? The Francis Scott Key Bridge on the Patapsco River.
What is the name of the ship that collided with the baltimore bridge? The ship was named the Dali.
How many articles did CNN publish about the Baltimore bridge collapse? CNN published 5 articles.
On what date did the Baltimore Bridge collapse? The collapse occurred on March 26, 2024.
Who is Brandon Scott? Brandon Scott is the Mayor of Baltimore.
Who is Niki Fennoy? Niki Fennoy is a city police spokesman.
What are the top themes present in the data? Themes include maritime incidents, manmade disaster, and water-related topics.
Table 2: Ground Truth Answers for the Baltimore Bridge Collapse Questions
To further validate our qualitative findings, Figure 5 presents a quantitative comparison of the semantic similarity of answers generated by the five different methods when querying the GDELT dataset related to the Baltimore bridge collapse. We compute the quality of the result by calculating cosine similarity [^5] The semantic embeddings used to calculate the cosine similarity were generated using the sentence-transformers/all-MiniLM-L6-v2 model from the Sentence Transformers library., a measure of the similarity between the embeddings of the predicted answer and the ground truth, with higher values indicating greater accuracy. The box plots illustrate the distribution of these similarity scores for each method: Graph Query on DKG, G-Retriever on DKG, RAG using Vector Store, G-Retriever on LKG, and GraphRAG Q&A on GRKG. As our qualitative analysis suggested, methods leveraging direct graph queries on the DKG, standard RAG with a vector store, and GraphRAG Q&A on the GRKG generally exhibit the highest cosine similarity scores, providing quantitative evidence for their superior performance in generating accurate and relevant answers compared to the G-Retriever, especially with the LKG. This visualization quantitatively confirms the trends observed in the qualitative evaluation presented in Table 1.
Figure 5: Box plots comparing the cosine similarity scores of different question-answering methods applied to the GDELT data concerning the Baltimore bridge collapse. Higher cosine similarity indicates a greater semantic similarity between the predicted and actual answers.